Collect like terms and simplify the expression:




12m2 – 9m + 5m – 4m2 – 7m + 10

1
Expert's answer
2021-11-23T17:16:29-0500

"12m^2 - 9m + 5m - 4m^2 – 7m + 10\\
=12m^2 - 4m^2– 9m + 5m - 7m + 10\\
=(12m^2 - 4m^2)+ (5m - 9m - 7m) + 10\\
=8m^2-11m+10"